OpenStudy (anonymous):

I need help on these. I know this is easy but im in a bad mood..please help me :( 2. –2x + 1 ≤ –11= 3. 10 < –3x + 1 4. 2(x + 5) > 8x – 8 5. –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

this would help me to. my creative juices aren't flowing at the moment Describe a real life scenario where inequalities are used and state the inequality.

OpenStudy (anonymous):

2. –2x + 1 ≤ –11= x≥6 3. 10 < –3x + 1= x<-3 4. 2(x + 5) > 8x – 8= x<3 5. –2(x – 3) ≥ 5 – (x + 3)= x≤4

OpenStudy (anonymous):

thankyou